OpenStudy (anonymous):

solution for system of equation for 3x+2y=4 5x-3y=3

OpenStudy (anonymous):

Alright, this is easy. For a system of equations, we want to get rid of one of the variables, so to do that we should make their coefficients the same. (For either x or y) For my example, I'll make y equal to each other. If we multiply 3 to the first equation to make it 9x+6y=12 and multiply 2 to the second one to make it 10x-6y=6, then we can 9x+6y=12 + 10x-6y=6 19x=18, and you can solve from there.